OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sin2x + sin x = 0

OpenStudy (anonymous):

\(\sin(2x)\) or \(\sin^2(x)\)?

OpenStudy (anonymous):

sin2x=2sinxcosx so 2sinxcosx+sinx=0 sinx(2cosx+1)=0 this will be 0 whne either sinx =0 or cosx=-1/2

OpenStudy (anonymous):

sorry, its sin^2(x)

OpenStudy (anonymous):

Find the roots of \(x^2+x=0\) and then solve for \(\sin(\theta)=x\)

OpenStudy (anonymous):

ohh, then make sinx=t t^2+t=0 t(t+1)=0 so t=0, or t=-1 substituting back sinx=0 or sinx=-1 means x=0 or x=3Pi/2

OpenStudy (anonymous):

@myko Is correct, but we mustn't forget that \(\sin(\pi)=0, \pi\in[0,2\pi)\)

OpenStudy (anonymous):

so my answers will be 0, pi, and 3pi/2 then?

OpenStudy (anonymous):

right. x=0 or Pi or x=3Pi/2

OpenStudy (anonymous):

thanks very mcuh!!